OpenStudy (anonymous):

The body mass index ,I, can be used to determine an individual's risk for heart disease. An index less than 25 indicates a low risk. The body mass index is given by the formula, or model, I=700W/H2 where W=weight in pounds and H= height in inches. Jerome is 77 inches tall. What weights will keep his body mass index between 18 and 26

OpenStudy (anonymous):

Is that H squared or 2H?

OpenStudy (anonymous):

If it's H^2, it makes a lot more sense. For the smallest weight: 18 = 700W/5929; W = 152.46 For the largest weight: 26 = 700W/5929; W = 220.22

OpenStudy (anonymous):

Ok it is h squared for Jerome that means Index = 700w/77^2 or Index = 700w/5929 multiply each side by 5929 5929I = 700w set index to one of the number 5929(18) = 700w 106722 = 700 w divide by 700 each side w = 152.46....and what danial said =)